Argon is contained in a 3 m
3
rigid volume at 75 C and 400 kPa. Calculate the heat transfer
needed to increase the pressure to 800 kPa.
1
Expert's answer
2019-02-26T06:55:42-0500

Relevant equations are:

Q = mCΔT = ρVC(T2 - T1),

P2/T2 = P1/T1

where

heat transferred: Q

density of argon: ρ = 1.784 kg/m³

specific heat: C = 0.52 kJ/kg K

initial temperature: T1 = 75+273 = 348 K

initial pressure: P1 = 400 kPa

final pressure: P2 = 800 kPa

T2 = (P2/P1)T1 = (800/400)348 = 696 K

Q = ρVC(T2 - T1) = 1.784*3*0.52*(696-348) = 968.5 J